One number is two more than a second number. Twice the first number is triple the second number minus 4. What are the numbers?
Answer by macston(5194)

.
F=first number; S=second number
.
F=S+2
.
2F=3S-4
F=(3/2)S-2
.
(3/2)S-2=S+2
(1/2)S=4
S=8
.
F=S+2=8+2=10
.
ANSWER: The first number is 10 and the second is 8.
.
CHECK:
2F=3S-4
2(10)=3(8)-4
20=24-4
20=20